the ratio of the two acute angles in a right triangle is 5:13. what is the measure of each angle in the right triangle?
Answer by lwsshak3(11628)

the ratio of the two acute angles in a right triangle is 5:13. what is the measure of each angle in the right triangle?
***
let x=measure of one acute angle
90-x=measure of other acute angle
x/(90-x)=5/13
13x=450-5x
18x=450
x=25
90-25=65
measure of one acute angle=25˚
measure of other acute angle=65˚